OpenStudy (anonymous):

Simplify: x^2-8(2x-8) / 64-x^2

OpenStudy (anonymous):

x^2-8(2x-8) / 64-x^2 x^2-16x+64 / 64-x^2 x^2-16x+64 = (x - 8)(x - 8) 64-x^2 = (8 - x)(8 + x) (x - 8)(x - 8) / (8 - x)(8 + x) do the rest

OpenStudy (anonymous):

can you do it or you need help